OpenStudy (anonymous):

tan x = sqrt(3) (unit circle problem)

OpenStudy (jdoe0001):

\(\bf tan(x)=\sqrt{3}\qquad \textit{taking inverse tangent to both sides}\\ \quad \\ tan^{-1}[tan(x)]=tan^{-1}[\sqrt{3}]\implies x = tan^{-1}[\sqrt{3}]\)

OpenStudy (anonymous):

but its by using the unit circle. the answer is (pi/3)+pi k, but i dont know how to get that answer

OpenStudy (anonymous):

Use your special triangles.

OpenStudy (anonymous):